OpenStudy (anonymous):

Find m(5) using direct substitution: m(x)=x+x^2-1 m(5)=(5)+(5)^2-1 How do you do direct substitution?

OpenStudy (anonymous):

@DemolisionWolf do you know what "direct substitution" is?

OpenStudy (anonymous):

cant really explain with this situation sorry.

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

@ali1029 this guy was getting mad at me for asking him to stop handing out answer while i'm trying to teach ^_^ haha direct substitution is what you show in the question post

OpenStudy (anonymous):

@ali1029 I dont see any other possible way you can do anytype of substitution aside from what you've done, sythetic or direct or whatever

OpenStudy (anonymous):

yeah @DemolisionWolf is right. you substituted 5 in for x, that's an example of direct substitution.

OpenStudy (anonymous):

haha well thats crazy! you're supposed to actually learn the material :p & okay, so even though it specifically says to use "direct substitution" I could just do m(5)=(5)+(5)^2-1 which would be 25+5-1=29?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

ya, as far as I can tell. i'm trying to dig up some facts on the differences between the two types

OpenStudy (anonymous):

sound good, let me know if there are any changes! :)

OpenStudy (anonymous):

thank you also @carrolldy! :)

OpenStudy (anonymous):

no problem. im good in math, if you need help hmu! :)

OpenStudy (anonymous):

Wow, thanks for offering that! I definitely will :p

OpenStudy (anonymous):

@ali1029 @carrolldy so here is the difference between the direct substitution and sythetic direct is like this: (what u did) m(x)=x+x^2-1 m(5)=(5)+(5)^2-1 synthetic is a long process of this: m(x)=x+x^2-1 5 | 1 1 <-pull down the coefficients in front of the x terms 5 25 <-do the terms x and x^2 so, 5 and 5^2 or 5 and 25 5 25 -1 <-then multiply the 1st row by the 2nd row and add the constants 29 <- then add up all the terms on the 3rd row to get the answer so... synthetic is a way i think to teach substitution, but clearly the 3 of us are more advanced than this and just do direct substitution

OpenStudy (anonymous):

ahhh, so my final answer of 29 is correct? we're so advanced ;)

OpenStudy (anonymous):

totes advanced ^_^

OpenStudy (anonymous):

hahaha :D